### v1

# Exciton fractional Chern insulators in moiré heterostructures

Abstract
Moiré materials have emerged as a powerful platform for exploring exotic quantum phases. While recent experiments have unveiled fractional Chern insulators exhibiting the fractional quantum anomalous Hall effect based on electrons or holes, the exploration of analogous many-body states with bosonic constituents remains largely uncharted. In this work, we predict the emergence of bosonic fractional Chern insulators arising from long-lived excitons in a moiré superlattice formed by twisted bilayer WSe2 stacked on monolayer MoSe2. Performing exact diagonalization on the exciton flat Chern band present in this structure, we establish the existence of Abelian and non-Abelian phases at band filling $\frac{1}{2}$ and $1$, respectively, through multiple robust signatures including ground-state degeneracy, spectral flow, many-body Chern number, and particle-cut entanglement spectrum. The obtained energy gap of $\sim 10$ meV for the Abelian states suggests a remarkably high stability of this phase. Our findings not only introduce a highly tunable and experimentally accessible platform for investigating bosonic fractional Chern insulators but also open a new pathway for realizing non-Abelian anyons.
I Introduction
Fractional Chern insulators (FCIs)—lattice analogs of the fractional quantum Hall effect that remain robust in the absence of a magnetic field—hold large potential for the study of fundamental quantum phenomena and for the development of novel quantum technologies [1, 2, 3]. Pioneering experiments [4, 5] and predictions [6, 7, 8] of FCIs in twisted van der Waals heterostructures, followed by recent realizations at absent magnetic field [9, 10, 11, 12], have established moiré materials as an accessible and versatile platform for exploring strongly-correlated topological phases. This breakthrough has stimulated abundant efforts in the pursuit of exotic phases beyond the conventional paradigm of Laughlin and hierarchy fractional quantum Hall states. In particular, predictions of Moore–Read [13, 14, 15, 16, 17, 18] and Read–Rezayi [19] phases hosting non-Abelian anyon excitations are especially promising for fault-tolerant topological quantum computing [20]. So far, however, research on moiré FCIs has focused only on the approach of doping the system with electrons or holes—resulting in correlated topological phases with fermionic constituents and leaving their bosonic counterpart largely unexplored.
Excitons, i.e. Coulomb-bound pairs of conduction-band electrons and valence-band holes [21], are obvious candidates for realizing correlated bosonic phases in moiré materials [22, 23]. Concretely, interlayer excitons with charge-carriers located in different layers are particularly promising due to their long lifetime, which can reach hundreds of nanoseconds [24, 25] and even microseconds [26]. These species typically appear in semiconducting van der Waals heterostructures, where an optical excitation generating intralayer excitons (with electrons and holes in the same layer) is shortly followed by tunneling of either electrons or holes into another layer [27]. The resulting interlayer excitons possess a permanent dipole moment, which makes them highly tunable by an out-of-plane electric field [28, 29]. Moreover, repulsive dipolar interactions together with the presence of flat bands in moiré semiconductors lead to correlated exciton physics  [30, 31, 32, 33]. In addition, the large degree of control via twist angle, dielectric environment, and electric field can be exploited to achieve topological bands [34] hosting long-lived interlayer excitons [35]. Thus, moiré systems enable the realization of topological exciton flat bands, providing a promising route for the exploration of exciton physics with intertwined correlations and topology.
Figure 1: (a) Schematic illustration of twisted bilayer WSe2 on top of monolayer MoSe2. The interlayer excitons X1 and X2 are formed by an electron in MoSe2 and a hole in either of the two WSe2 layers. (b) Exciton band structure for $\Delta=-3.8$ meV and $\theta=1.95^{\circ}$, where the lowest band is flat and has a non-zero Chern number. The color represents the contribution $|\chi_{l\mathbf{Q}}|^{2}$ from each exciton species to the band. (c) Berry curvature $\Omega_{\mathbf{k}}A_{\text{BZ}}/2\pi$ and (d) Fubini-Study metric $\text{tr}[g_{\mathbf{k}}]A_{\text{BZ}}/2\pi$ in the moiré Brillouin zone for the flat band. $A_{\text{BZ}}$ is the Brillouin zone area.
In this work, we unveil the many-body topological exciton phases emerging in a moiré heterostructure. First we show that, besides being topologically non-trivial, the lowest exciton band in twisted bilayer WSe2 stacked on monolayer MoSe2 can simultaneously be nearly flat and exhibit an almost ideal quantum geometry—posing this system as a strong candidate for realizing exciton FCI phases. Employing exact diagonalization, we show that the ground state at half filling with contact interactions is analogous to the bosonic Laughlin state in the lowest Landau level. Concretely, the state is characterized by a twofold degeneracy as well as an approximately zero (interaction) energy, and its nature is further confirmed by the many-body Chern number, spectral flow, and entanglement spectrum. Importantly, the Laughlin states remain robust and exhibit a large gap ($\sim 10$ meV) when replacing the idealized contact potential by realistic long-range interactions. Finally, the many-body calculations at filling $1$, again considering long-range interactions, provide compelling evidence for a stable bosonic version of the non-Abelian Moore–Read state. Overall, our findings predict the existence of exciton FCIs in an experimentally accessible moiré heterostructure and pave the way for the realization of Abelian and non-Abelian topological phases with bosonic constituents in moiré materials.
II Model
II.1 Topological exciton flat band
We consider a van der Waals heterostructure consisting of a twisted WSe2 bilayer stacked on top of a single MoSe2 layer. An optical excitation and subsequent tunneling result in the formation of long-lived interlayer excitons composed of a hole in either of the two WSe2 layers and an electron in MoSe2, cf. Fig. 1(a). We assume spin-valley polarization of electrons and holes, which can be achieved by a circularly-polarized optical excitation if the valley decoherence time is sufficiently long and, in some cases, occurs spontaneously due to interactions. In order to describe moiré excitons in this structure, we follow the procedure introduced in Ref. [35].
First, we set up the exciton basis, |$\ket{X_{l\mathbf{Q}}}=\sum_{\mathbf{k}}\phi_{l\mathbf{k}}e^{\dagger}_{\mathbf{k}_{\text{e}}}h^{\dagger}_{l\mathbf{k}_{\text{h}}}\ket{0}$, which accounts for the binding of an MoSe2 electron to a WSe2 hole in the layer $l=1,2$. Details of the exciton basis, including the relation between the electron/hole momentum $\mathbf{k}_{\text{e/h}}$ and the relative and center-of-mass momenta $\mathbf{k}$ and $\mathbf{Q}$ can be found in Appendix A.1. The exciton’s wave function $\phi_{l\mathbf{k}}$ and binding energy $E^{\text{x}}_{l}$ are obtained by solving the Schrödinger equation for an electron–hole pair interacting via the interlayer 2D Coulomb potential. Here, we describe electrons and holes in an effective-mass approximation, where the exciton eigenvalue problem takes the form of the Wannier equation. In this case, $\phi_{l\mathbf{k}}$ is well described by the 1s wave function of 2D hydrogen, and variational minimization considering material-specific parameters yields the exciton binding energies $E^{\text{x}}_{1}=230\ \text{meV},E^{\text{x}}_{2}=140\ \text{meV}$ in close agreement with the exact solutions.
Next, we incorporate the effect of the moiré potential and tunneling that holes in the WSe2 layers experience, which is described by the moiré exciton Hamiltonian [35, 36]
$H_{\text{x,m}}=\sum_{ll^{\prime}\mathbf{Qq}}M_{ll^{\prime}\mathbf{Qq}}X^{\dagger}_{l\mathbf{Q+q}}X^{\phantom{\dagger}}_{l^{\prime}\mathbf{Q}},$ (1)
where mixing with higher exciton states (e.g. 2p) is expected to be weak and has therefore been disregarded. We note that this approach naturally accounts for the mixing between all moiré hole bands forming the exciton state. Here, $M_{ll\mathbf{Qq}}=E^{\text{x}}_{l\mathbf{Q}}\delta_{\mathbf{q},0}+\mathcal{F}_{ll,\alpha_{\text{e}}\mathbf{q}}U_{\mathbf{q}}$ contains the exciton dispersion $E^{\text{x}}_{l\mathbf{Q}}=E^{\text{gap}}_{l}+E^{\text{x}}_{l}+\hbar^{2}\mathbf{Q}^{2}/2M_{\text{x}}$ with the exciton mass $M_{\text{x}}$, the hole moiré potential $U_{\mathbf{q}}=U_{0}\sum_{n=1}^{6}\mathrm{e}^{i\sigma_{nl}\varphi}\delta_{\mathbf{q},\mathbf{g}_{n}}$ with $\sigma_{nl}=(-1)^{n+l-1}$, $\mathbf{g}_{n}=C^{n}_{6}\mathbf{g}_{0}$, $\mathbf{g}_{0}=4\pi/\sqrt{3}a_{\text{m}}\mathbf{\hat{x}}$, the moiré lattice constant $a_{\text{m}}$ and the exciton form factor $\mathcal{F}_{ll^{\prime},\mathbf{q}}=\sum_{\mathbf{k}}\phi^{*}_{l\mathbf{k}}\phi_{l^{\prime}\mathbf{k+q}}$. Tunneling between opposite layers is contained in $M_{ll^{\prime}\mathbf{Qq}}=\mathcal{F}_{ll^{\prime},\alpha_{\text{e}}\mathbf{q}}T_{\mathbf{q}}$ ($l\neq l^{\prime}$), where $T_{\mathbf{q}}=T_{0}\sum_{n=1}^{3}\delta_{\mathbf{q},{\bm{\kappa}}_{n}}$, ${\bm{\kappa}}_{n}=C^{n}_{3}(\mathbf{K}_{2}-\mathbf{K}_{1})$ and $\mathbf{K}_{l}$ is the K-point in layer $l$. The relevant parameters for holes in twisted bilayer WSe2 are $\varphi=128^{\circ}$, $U_{0}=9$ meV, and $T_{0}=18$ meV [35]. Diagonalizing Eq. (1) yields the band structure, which in this system is characterized by topological bands for a specific range of the twist angle and the offset between the two exciton species, $\Delta=E^{\text{x}}_{2}-E^{\text{x}}_{1}$. The latter can be experimentally controlled by an out-of-plane electric field.
In this work, we consider the twist angle $\theta=1.95^{\circ}$ and the offset $\Delta=E^{\text{x}}_{2}-E^{\text{x}}_{1}=3.8$ meV, where the lowest band is flat and has a Chern number $C=1$, cf. Fig. 1(b). The non-zero Chern number can be understood in terms of a pseudospin, describing the superposition between the two interlayer exciton species |$\ket{X_{1\mathbf{Q}}}$ and |$\ket{X_{2\mathbf{Q}}}$, which wraps around the Bloch sphere once as $\mathbf{Q}$ traverses the moiré Brillouin zone. The varying pseudospin is encoded in the color of the flat band in Fig. 1(b), where the weight of the state |$\ket{X_{l\mathbf{Q}}}$, i.e. $|\chi_{l\mathbf{Q}}|^{2}$ with $\chi_{l\mathbf{Q}}$ being the moiré exciton eigenfunction, is shown. Note that $\chi_{l\mathbf{Q}}$ is a vector where each element accounts for the momentum $\mathbf{Q}+\mathbf{g}$ with $\mathbf{g}$ pointing to the outer moiré cells up to a certain cutoff. Interestingly, the quantum geometry of the band is nearly ideal [37, 38], i.e. $\mathrm{tr}[g_{\mathbf{k}}]\approx|\Omega_{\mathbf{k}}|$ where §$\textsection g_{\mathbf{k}}$ is the quantum (Fubini–Study) metric and $\Omega_{\mathbf{k}}$ is the Berry curvature, cf. Fig. 1 (c), (d). This property strongly suggests the emergence of zero-energy ground states at even-denominator filling of the band with pseudopotential interactions—in analogy to Laughlin states of bosons in Landau levels, albeit now in the absence of a magnetic field. A nearly ideal quantum geometry was also found to be the precursor of electron FCIs in twisted bilayer graphene [8, 39] despite the fact that fluctuations of the metric induce new competing states [40].
II.2 Many-body exciton Hamiltonian
The aim of this work is to unveil the phases emerging in a many-body system of excitons in the topological moiré flat band. To that end, we set up the many-body exciton Hamiltonian, $H_{\text{x}}=H_{\text{x,m}}+H_{\text{x-x}}$, where
$H_{\text{x-x}}=\frac{1}{2}\sum_{ll^{\prime}\mathbf{q}}V_{ll^{\prime}}(\mathbf{q}):\!\rho_{l}(\mathbf{q})\rho_{l^{\prime}}(-\mathbf{q})\!:$ (2)
describes the exciton-exciton interaction, $\rho_{l}(\mathbf{q})=\sum_{\mathbf{Q}}X^{\dagger}_{l\mathbf{Q+q}}X^{\phantom{\dagger}}_{l\mathbf{Q}}$ is the exciton density operator, :::absent:::: : denotes normal order, and
$V_{ll^{\prime}}(\mathbf{q})=$ $V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})+V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})-V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})-[V^{\text{e-h}}_{l^{\prime}l}(-\mathbf{q})]^{*}$
$V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{h}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}$
$V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{e}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|d|l-l^{\prime}|}$
$V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|dl}$ (3)
is the interaction potential, which contains individual contributions accounting for interactions between the charges in different excitons, and where $F_{ll^{\prime}}(\mathbf{q}_{1},\mathbf{q}_{2})=\mathcal{F}_{ll,\mathbf{q}_{1}}\mathcal{F}_{l^{\prime}l^{\prime},-\mathbf{q}_{2}}$ has been introduced. The derivation of the exciton–exciton interaction potential is outlined in Appendix A.2. Here, we have focused on the direct interaction, which at long distances behaves as the interaction between two dipoles with lengths $dl$ and $dl^{\prime}$, i.e. $V_{ll^{\prime}}(\mathbf{r})\sim(dl)(dl^{\prime})/|\mathbf{r}|^{3}$. Exchange interactions typically play a minor role in the context of interlayer excitons [41, 42, 43, 44].
Importantly, the exciton operators $X^{\dagger}_{\mathbf{Q}}$ and $X^{\phantom{\dagger}}_{\mathbf{Q}}$ obey bosonic commutation relations. Corrections to the bosonic character arising from the exciton’s fermionic substructure are typically expected in the regime where the distance between neighboring excitons is comparable to the size of an exciton [45], i.e. $n_{\text{x}}a_{\text{B}}^{2}\sim 1$ where $n_{\text{x}}$ is the exciton density and $a_{\text{B}}$ the exciton Bohr radius. In typical moiré systems where there is one exciton per moiré site, $n_{\text{x}}a_{\text{B}}^{2}\sim 0.01$ and thus the bosonic description is appropriate.
The many-body problem of interacting excitons is solved via exact diagonalization. Concretely, we project the exciton Hamiltonian $H_{\text{x}}$ into the topological flat band and we diagonalize $H_{\text{x}}$ in a finite-size system consisting of $N_{\text{x}}$ excitons in $N_{\text{s}}$ moiré sites such that the band filling is $\nu=N_{\text{x}}/N_{\text{s}}$. We distinguish exciton FCIs from competing orders by assessing the ground state degeneracy, many-body Chern number (defined in Appendix B.1), and quasi-hole excitations. The latter can be accessed via the particle-cut entanglement spectrum (PES) [46, 47], where the many-body system is partitioned into $N_{A}$ and $N_{B}=N_{\text{x}}-N_{A}$ particles. The set of eigenvalues $\{\xi\}$ of $-\mathrm{log}\rho_{A}$ consitute the PES, where $\rho_{A}=\text{tr}_{B}[\frac{1}{N_{d}}\sum_{i}^{N_{d}}|\Psi_{i}\rangle\langle\Psi_{i}|]$ is the reduced density matrix of the subsystem $A$, $|\Psi_{i}\rangle$ is the $i$-th ground state, and $N_{d}$ is the ground state degeneracy. Writing the many-body ground state as |$\ket{\Psi}=\sum_{j}\mathrm{exp}(-\xi_{j}/2)\ket{\Psi^{A}_{j}}\otimes\bra{\Psi^{B}_{j}}$ |, one can see that lower values of $\xi_{j}$ indicate a large weight of a specific subspace configuration |$\ket{\Psi^{A}_{j}}$—meaning that such configuration is allowed. Thus, the PES probes quasi-hole excitations that obey the generalized statistics of a specific quantum phase.
Figure 2: Laughlin states at half filling. (a) Many-body energy spectrum containing the 10 lowest energies for each momentum sector, (b) spectral flow, (c) particle-cut entanglement spectrum, and (d) many-body Berry curvature for the two ground states considering contact interactions without kinetic energy effects at $\nu=\frac{1}{2}$. The respective data considering the long-range interaction $V_{ll^{\prime}}(\mathbf{q})$ and the kinetic energy is shown in (e)-(h). The ground states in (a) and (e) are marked in red and have an average many-body Chern number $C_{\text{avg}}=1/2$. In the PES, the number of states below the first entanglement gap (denoted by the red solid line) is $1287$, matching the number of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states. The considered system has $N_{\text{s}}=18$ moiré sites and its spanning vector can be found in the Appendix B.2
III Abelian states at half filling
Based on the criteria for how closely a moiré band can mimic a Landau Level—(i) nontrivial topology, (ii) flat dispersion, and (iii) ideal quantum geometry [37, 3, 38]—the exciton flat band shown in Fig. 1(b) appears to be an excellent candidate for realizing FCI phases analogous to the fractional quantum Hall effect of bosons. In this context, Laughlin states at half filling constitute a prototypical phase [48, 49]. They emerge as exact zero-energy states in a system of bosons interacting via a delta-function potential (i.e. contact interaction) in the lowest Landau level and their elementary excitations behave as Abelian anyons. We note that bosonic Laughlin states have also been proposed in a different setting, where the exciton-like bosons correspond to low-energy excitations of a fully-filled valley-polarized electron band [50, 51, 52]. In the following, we provide numerical evidence showing that the nearly ideal exciton Chern band indeed supports zero-energy Laughlin states for contact interactions. Subsequently, we will demonstrate that this FCI phase survives when considering the realistic long-range interactions introduced in Eq. (3).
In order to strengthen the ideal conditions of the system and establish the emergence of Laughlin states in the exciton Chern band, we initially consider contact interactions $V_{ll^{\prime}}(\mathbf{q})=V_{0}$ and disregard the impact of the kinetic energy. With these assumptions, exact diagonalization yields two many-body ground states with approximately zero energy, cf. Fig. 2(a). Importantly, the ground states are separated from excited states by a large energy gap, which reflects the expected topological protection. The twofold degeneracy and the total momentum of each state are characteristic of Laughlin states and can be understood with the aid of Landau level physics in the thin-torus limit [53]. In particular, the exclusion principle that arises in the thin-torus limit for bosons in the lowest Landau level dictates that a zero-energy state contains, at most, one particle in two consecutive sites [54]. The two states that fulfill this principle and therefore constitute the two degenerate ground states are those with a Fock-space configuration $101010\cdots$ and its translational-invariant partner $010101\cdots$, whose total momentum matches that of our numerical ground states. Furthermore, upon threading a magnetic flux (corresponding to twisted boundary conditions), the ground states evolve into each other after one flux quantum and return to their original states after the insertion of two flux quanta (Fig. 2(b)), indicating that the two states are adiabatically connected and that the Hall conductivity is quantized to $\frac{1}{2}$. The latter aspect is further confirmed by a direct calculation of the many-body Chern number, which yields $C_{\text{avg}}=\frac{1}{2}$ for each ground state and arises from a homogeneous many-body Berry curvature, cf. Fig. 2(d).
Next, we calculate and analyze the PES, which reveals the nature of the elementary quasi-hole excitations of ground states. In particular, the states with lowest eigenvalues in the PES typically correspond to hole excitations that fulfill the generalized exclusion principle and therefore serve as a fingerprint for distinguishing a specific phase from competing orders. In Fig. 2(c), we display the PES where the subsystem $A$ is taken to consist of $N_{A}=4$ particles. Here, states with low eigenvalues are well separated from higher eigenvalues by a large entanglement gap. Importantly, the number of states below the gap exactly matches the analytical counting of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states. Thus, the many-body ground state degeneracy, spectral flow, many-body Chern number, and PES all taken together clearly confirm the emergence of a stable FCI phase of excitons in the half-filled moiré Chern band.
We now consider the realistic long-range exciton–exciton interactions introduced in Eq. (3) and take into account the impact of the small but finite kinetic energy. The results of exact diagonalization and the subsequent analysis are remarkably similar to those obtained for contact interactions, cf. Fig. 2(e)-(h). The fact that the energy gap between ground and excited many-body states ($\sim 10$ meV) persists for the considered systems ranging from $N_{\text{s}}=10$ up to $N_{\text{s}}=20$ sites clearly reflects that such phases are robust (see Appendix C.1). Thus, our numerical calculations demonstrate that exciton FCIs are stable in the considered moiré heterostructure.
Interestingly, the calculated gap of $\sim 10$ meV is two times larger than in the case of hole FCIs in the similar system of twisted bilayer MoTe2 [55]. In principle, this suggests a higher stability of exciton FCIs compared to their electron/hole counterparts, whose gap has been experimentally estimated to be 20 K [56]. The potentially larger gap of exciton FCIs might be a result of the stronger repulsive interaction dominated by $V^{\text{e-e}}+V^{\text{h-h}}$ in the short range ($V^{\text{e-h}}$ is weaker since the charges are located in separate layers) instead of just $V^{\text{e-e}}$ or $V^{\text{h-h}}$. A larger gap for bosons can also be expected from a pseudopotential perspective, where the interaction for bosons and fermions is dominated by the zeroth, $v_{0}$ ($\delta(\mathbf{r)}$ component of the interaction) and first $v_{1}$ ($-a^{2}_{\text{m}}\nabla^{2}\delta(\mathbf{r)}$ component) pseudopotentials, respectively, with the bosonic interaction being stronger than the fermionic one (generally $v_{n}>v_{n+1}$ in lowest Landau level-like bands) [49, 8]. We note, though, that a reliable quantitative estimation of the gap is difficult due to limitations such as the finite system size, the use of a pure Coulomb potential instead of the thin-film Rytova-Keldysh potential [21], and uncertainties in parameters such as the dielectric constant.
IV Non-Abelian states at filling one
After confirming the presence of Abelian exciton FCIs, we now seek exotic phases whose elementary excitations obey non-Abelian anyon statistics. In this context, the most straightforward and promising candidate is the Moore–Read state [57], which emerges as the exact zero-energy ground state at filling $\nu=1$ for bosons in the lowest Landau level with artificial three-body contact interactions [58] and which was predicted to appear in rotating Bose-Einstein condensates [59, 60]. In the following, we show that this phase is stable in the realistic conditions considered here, i.e. in the moiré band and assuming long-range interactions.
Figure 3: Moore–Read states at filling one. Many-body spectrum (10 lowest energies) and PES at $\nu=1$ with long-range interactions for (a)-(b) 13 and (c)-(d) 14 particles. The single and threefold quasi-degenerate ground states (red dots in panels (a) and (c)) for an odd and even number of sites are located at the momenta expected for Moore–Read states. The red line in panels (b) and (d) indicates the entanglement gap respecting the quasi-hole counting rules of the Moore–Read state (416 and 518 quasihole excitations for 13 and 14 particles with $N_{A}=3$). PES for larger $N_{A}$ are shown in Appendix C.2.
First, we note that the most prominent feature of the Moore–Read phase is the dependence of the ground state degeneracy on the parity of the particle number. Here, the thin-torus exclusion principle [61] states that, at most, two particles can occupy two consecutive orbitals [62, 54], resulting in three possible Fock-space configurations, $111111\cdots$, $202020\cdots$, and $020202\cdots$. As a result of periodic boundary conditions, only the first configuration is allowed for an odd number of particles—resulting in a single ground state—, while for an even number of particles all three configurations are allowed and the ground state is threefold degenerate.
In Fig. 3(a),(c), we show the calculated many-body spectra for systems with 13 and 14 excitons. The single and three lowest states in the case of odd and even number of excitons, respectively, are located at the momenta expected from the thin-torus exclusion rule for Moore–Read states. We note that the spread of the ground state energies and the small gap with respect to excited states is a common feature of finite-size systems. We expect the states to become exactly degenerate in the thermodynamic limit and note that the ground and excited states remain well separated upon threading magnetic flux (see Appendix C.2). Moreover, the calculation of many-body Chern number yields $\mathcal{C}_{\text{avg}}=1$ for each state and reveals a uniform Berry curvature distribution, indicating that the Moore–Read phase is robust.
In order to unambiguously determine the nature of the phase beyond the parity dependence of the ground-state degeneracy, we calculate the PES. For both $N_{\text{x}}=13$ and $N_{\text{x}}=14$ systems, the PES is characterized by a large gap, below which the number of states exactly matches the counting of quasi-hole excitations allowed by the Moore–Read exclusion rule. Furthermore, the PES gap persists across various system sizes and different $N_{A}$ (see Appendix C.2). Taken together, the ground-state degeneracy and momenta, the many-body Chern number, and the PES for systems with different particle-number parity constitute convincing evidence for the presence of a robust non-Abelian Moore–Read phase with exciton constituents at filling $\nu=1$ in the considered moiré heterostructure.
V Conclusion and outlook
We have explored the emergence of strongly-correlated topological phases arising from a many-body bosonic system of long-lived moiré excitons in a flat Chern band. In particular, by combining exact diagonalization with many-body diagnosis tools, we have demonstrated the existence of robust Abelian and non-Abelian exciton fractional Chern insulators in a realistic model of a moiré heterostructure at filling $\nu=\frac{1}{2}$ and $\nu=1$, respectively. Our work introduces a versatile and accessible platform for investigating exciton FCIs and opens a new avenue towards the realization of non-Abelian anyons.
Importantly, our findings provide a specific guideline for the experimental realization of exciton FCIs in twisted bilayer WSe2 stacked on monolayer MoSe2. We note that, despite the exciton’s neutral net charge, there are methods which enable exciton transport that could be utilized to detect these phases. The most promising route is likely to be counterflow transport measurements [63, 64]. Other methods for studying transport of interlayer excitons involve creating spatial gradients in the out-of-plane electric field, dielectric environment, or strain profiles [65, 66]. Future studies should also attempt to identify the optical fingerprints of these phases.
Apart from prompting experimental efforts in a specific system, our work motivates the search for exciton FCIs in other moiré structures. Furthermore, while we have only focused on two specific phases at fillings $\nu=\frac{1}{2}$ and $\nu=1$, other highly nontrivial phases remain unexplored. Besides other bosonic Laughlin and hierarchy states, the pursuit of additional non-Abelian phases such as the Moore–Read state at $\nu=\frac{1}{3}$ and the Read–Rezayi state [67] at $\nu=\frac{3}{2}$ is particularly interesting—especially so for the latter as it hosts Fibonacci anyons which hold large promise for topological quantum computing.
Finally, we note that many challenging aspects are yet to be addressed. For instance, investigating the competition between liquid- and crystal-like exciton FCIs as well as superfluids and supersolids would be crucial for the understanding of these phases [68, 69]. Such phase transitions could be studied by adding hBN spacers between the electron layer and the hole bilayer to enhance the interaction strength (by reducing the electron–hole attraction). In addition, while the spin-valley physics in the many-body exciton system is at the edge of current numerical capabilities, it is an important issue that must be tackled. Moreover, theoretical efforts exploring large filling factors might demand more advanced models considering deviations from the bosonic description of excitons [70]. Last but not least, the phase diagram can be enriched by adding doping, which in combination with excitons constitutes a realization of mixed Bose–Fermi physics [71, 33, 72]. Our work shows that an idealized model can provide a faithful description of moiré exciton FCIs, suggesting a route to approach these formidable problems.
Acknowledgements.
We acknowledge useful discussions with Daniel Erkensten, Zhao Liu, Ahmed Abouelkomsan and Ataç İmamoğlu. This work was supported by the Swedish Research Council (VR, grant 2024-04567), the Wallenberg Scholars program of the Knut and Alice Wallenberg Foundation (2023.0256) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. In addition, we utilized the Sunrise HPC facility supported by the Technical Division of the Department of Physics, Stockholm University.
Appendix A Moiré exciton model
A.1 Exciton basis
Electrons and holes in each layer can be described in an effective mass approach at the K valley with the Hamiltonian $H_{\text{e-h}}=H_{\text{e},0}+H_{\text{h},0}+H_{\text{e-h, int}}$, where $H_{\text{e},0}=\sum_{\mathbf{k}}E^{\text{e}}_{\mathbf{k}}e^{\dagger}_{\mathbf{k}}e^{\phantom{\dagger}}_{\mathbf{k}}$, $H_{\text{h},0}=\sum_{l\mathbf{k}}E^{\text{h}}_{\mathbf{k}}h^{\dagger}_{l\mathbf{k}}h^{\phantom{\dagger}}_{l\mathbf{k}}$, and $H_{\text{e-h, int}}=-\sum_{l\mathbf{kk^{\prime}q}}V^{\text{e-h}}_{l\mathbf{q}}e^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{l\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{l\mathbf{k}^{\prime}}e^{\phantom{\dagger}}_{\mathbf{k}}$. Here, $E^{\text{e}}_{\mathbf{k}}=\hbar^{2}\mathbf{k}^{2}/2m_{\text{e}}$, $E^{\text{h}}_{\mathbf{k}}=E^{\text{gap}}_{l}+\hbar^{2}(\mathbf{k}-\mathbf{K}_{l})^{2}/2m_{\text{h}}$, $\mathbf{K}_{l}$ is the K-point for holes at the layer $l=1,2$ and $V^{\text{e-h}}_{l\mathbf{q}}=\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|dl}$ is the electron-hole interaction potential with the interlayer distance $d$, the dielectric constant $\epsilon$, and the system area $A$. Considering that all electrons and holes are paired into tightly bound excitons, the system can be described in terms of the 1s exciton state, |$\ket{X_{l\mathbf{Q}}}=\sum_{\mathbf{k}}\phi_{l\mathbf{k}}e^{\dagger}_{\mathbf{k}_{\text{e}}}h^{\dagger}_{\mathbf{k}_{\text{h}}}\ket{0}$, where |$\ket{0}$ is the semiconductor ground state, $\phi_{l\mathbf{k}}$ is the exciton wave function with the relative momentum $\mathbf{k}=\alpha_{\text{h}}\mathbf{k}_{\text{e}}-\alpha_{\text{e}}(\mathbf{k}_{\text{h}}-\mathbf{K}_{l})$, $\alpha_{\text{e/h}}=m_{\text{e/h}}/(m_{\text{h}}+m_{\text{e}})$, and $\mathbf{Q}=\mathbf{k}_{\text{h}}-\mathbf{K}_{l}+\mathbf{k}_{\text{e}}$ is the center-of-mass momentum. The exciton dispersion $E^{\text{x}}_{l\mathbf{Q}}=E^{\text{gap}}_{l}+E^{\text{x}}_{l}+\frac{\hbar^{2}\mathbf{Q}^{2}}{2M_{\text{x}}}$ is determined by the mass $M_{\text{x}}=m_{\text{h}}+m_{\text{e}}$ and the offset energy $E^{\text{x}}_{l}$ that fulfills the eigenvalue problem $H_{\text{e-h}}\ket{X_{l\mathbf{Q}}}=E^{\text{x}}_{l\mathbf{Q}}\ket{X_{l\mathbf{Q}}}$, which corresponds to the Wannier equation
$\frac{\hbar^{2}\mathbf{k}^{2}}{2\mu_{\text{x}}}\phi_{l\mathbf{k}}-\sum_{\mathbf{q}}V^{\text{e-h}}_{l\mathbf{q}}\phi_{l,\mathbf{k+q}}=E^{\text{x}}_{l}\phi_{l\mathbf{k}},$ (4)
where $\mu_{\text{x}}=m_{\text{e}}m_{\text{h}}/M_{\text{x}}$ is the reduced mass. The exciton wave function is well approximated by a hydrogen ansatz, which in real space reads $\phi_{l}(\mathbf{r})\propto\mathrm{exp}(-|\mathbf{r}|/a_{l})$. Taking $m_{\text{h}}=0.43m_{0},m_{\text{e}}=0.8m_{0}$, $d=0.67$ nm and $\epsilon=3.8$[35] we obtain the exciton binding energies $E^{\text{b}}_{1}=230\ \text{meV},E^{\text{b}}_{2}=140\ \text{meV}$ and Bohr radii $a_{1}=1.3\ \text{nm},a_{2}=1.8\ \text{nm}$ via variational minimization, in close agreement with the exact numerical solution of Eq. (4).
A.2 Exciton–exciton interaction
We extend the method employed to derive the single-particle moiré exciton Hamiltonian [35] to obtain the Hamiltonian describing exciton–exciton interactions, which takes the form
$H_{\text{x-x}}=$ $\frac{1}{2}\sum_{\mathbf{QQ^{\prime}q}}V(\mathbf{q})X^{\dagger}_{\mathbf{Q+q}}X^{\dagger}_{\mathbf{Q^{\prime}-q}}X^{\phantom{\dagger}}_{\mathbf{Q^{\prime}}}X^{\phantom{\dagger}}_{\mathbf{Q}}$
$V(\mathbf{q})=$ $\braket{X_{\mathbf{Q+q}}X_{\mathbf{Q^{\prime}-q}}}{H_{\text{int}}}{X_{\mathbf{Q^{\prime}}}X_{\mathbf{Q}}}.$
For the sake of simplicity, we omit the layer indices and note that the generalization to the multilayer system is straightforward. The matrix element $V(\mathbf{q})$ contains the interaction Hamiltonian in the electron–hole picture,
$H_{\text{int}}=$ $H_{\text{e-e, int}}+H_{\text{h-h, int}}+H_{\text{e-h, int}}$
$H_{\text{e-e, int}}=$ $\frac{1}{2}\sum_{\mathbf{kk^{\prime}q}}V^{\text{e-e}}(\mathbf{q})e^{\dagger}_{\mathbf{k+q}}e^{\dagger}_{\mathbf{k^{\prime}-q}}e^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}e^{\phantom{\dagger}}_{\mathbf{k}}$
$H_{\text{h-h, int}}=$ $\frac{1}{2}\sum_{\mathbf{kk^{\prime}q}}V^{\text{h-h}}(\mathbf{q})h^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}h^{\phantom{\dagger}}_{\mathbf{k}}$
$H_{\text{e-h, int}}=$ $-\sum_{\mathbf{kk^{\prime}q}}V^{\text{e-h}}(\mathbf{q})e^{\dagger}_{\mathbf{k+q}}h^{\dagger}_{\mathbf{k^{\prime}-q}}h^{\phantom{\dagger}}_{\mathbf{k^{\prime}}}e^{\phantom{\dagger}}_{\mathbf{k}},$
The matrix element $V(\mathbf{q})$ thus has contributions from different interaction mechanisms. Here, we focus on the contribution arising from electron–electron interactions, $H_{\text{e-e, int}}$, to illustrate the procedure. First, we expand the exciton states into the electron–hole basis and exploit the fermionic commutation relations together with $e_{\mathbf{k}}\ket{0}=h_{\mathbf{k}}\ket{0}=0$ until all electron and hole operators in the matrix element disappear, obtaining
$\left.V(\mathbf{q})\right|_{\text{e-e}}=\sum_{\mathbf{k}_{i}}\phi^{*}_{\mathbf{k}_{1}}\phi^{*}_{\mathbf{k}_{2}}\phi_{\mathbf{k}_{3}}\phi_{\mathbf{k}_{4}}(A-B)C$
$A=\delta_{-\mathbf{k}_{2}-\alpha_{\text{h}}\mathbf{q},-\mathbf{k}_{3}}\delta_{-\mathbf{k}_{1}+\alpha_{\text{h}}\mathbf{q},-\mathbf{k}_{4}}$
$B=\delta_{-\mathbf{k}_{2}+\alpha_{\text{h}}(\mathbf{Q^{\prime}}-\mathbf{q}),-\mathbf{k}_{4}+\alpha_{\text{h}}\mathbf{Q}}\delta_{-\mathbf{k}_{1}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}),-\mathbf{k}_{3}+\alpha_{\text{h}}\mathbf{Q}^{\prime}}$
$C=\braket{0}{e^{\phantom{\dagger}}_{\mathbf{k}_{1}+\alpha_{\text{e}}(\mathbf{Q}+\mathbf{q})}e^{\phantom{\dagger}}_{\mathbf{k}_{2}+\alpha_{\text{e}}(\mathbf{Q}^{\prime}-\mathbf{q})}H_{\text{e-e}}e^{\dagger}_{\mathbf{k}_{3}+\alpha_{\text{e}}\mathbf{Q}^{\prime}}e^{\dagger}_{\mathbf{k}_{4}+\alpha_{\text{e}}\mathbf{Q}}}{0},$
where $i=\{1,2,3,4\}$. Next, we calculate $C$, again by making use of $e_{\mathbf{k}}\ket{0}=h_{\mathbf{k}}\ket{0}=0$ and the commutation relations, and evaluating the Kronecker deltas. The resulting expression reads
$\left.V(\mathbf{q})\right|_{\text{e-e}}=$ $\left.V(\mathbf{q})\right|_{\text{e-e,1}}-\left.V(\mathbf{q})\right|_{\text{e-e,2}}-\left.V(\mathbf{q})\right|_{\text{e-e,3}}+\left.V(\mathbf{q})\right|_{\text{e-e,4}}$
$\left.V(\mathbf{q})\right|_{\text{e-e,1}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}\mathbf{q}}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}\mathbf{q}}V^{\text{e-e}}(\mathbf{q})$
$\left.V(\mathbf{q})\right|_{\text{e-e,2}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}\mathbf{q}}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}\mathbf{q}}$
×$\times V^{\text{e-e}}(\mathbf{k}+\mathbf{q}-\mathbf{k}^{\prime}+\alpha_{\text{e}}(\mathbf{Q}-\mathbf{Q}^{\prime}))$
$\left.V(\mathbf{q})\right|_{\text{e-e,3}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}$
×$\times V^{\text{e-e}}(\mathbf{k}+\mathbf{q}-\mathbf{k}^{\prime}+\alpha_{\text{h}}(\mathbf{Q}-\mathbf{Q}^{\prime}))$
$\left.V(\mathbf{q})\right|_{\text{e-e,4}}=$ $\sum_{\mathbf{kk^{\prime}}}\phi_{\mathbf{k}}\phi_{\mathbf{k^{\prime}}}\phi^{*}_{\mathbf{k}+\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}\phi^{*}_{\mathbf{k}^{\prime}-\alpha_{\text{h}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime})}$
×$\times V^{\text{e-e}}(\mathbf{Q}+\mathbf{q}-\mathbf{Q}^{\prime}),$
where each term accounts for direct interaction, electron–electron exchange, hole–hole exchange, and exciton–exciton exchange, respectively [43]. For interlayer excitons, the direct term $V(\mathbf{q})|_{\text{e-e,1}}$ dominates [41] and therefore we disregard the remaining terms. Applying this method to the hole–hole and electron–hole interaction terms yields the exciton–exciton interaction potential in Eq. (3).
Appendix B Additional details of numerical methods
B.1 Many-body Chern number
In order to further assess the topological nature of ground states, we calculate the many-body Chern number, which takes the form
[…]
with $(\Phi_{x},\Phi_{y})$ and |$\ket{\Psi_{i}}$ being the inserted magnetic flux and the many-body ground state $i$, respectively. Here, the integral is evaluated over the $2\pi\times 2\pi$ torus ($T^{2}$). We calculate the Chern number for all many-body ground states and obtain the average many-body Chern number $\mathcal{C}_{\text{avg}}$ for each state [73].
B.2 Finite system geometry
Throughout this work, we consider finite systems (with periodic boundary conditions), the geometry of which is determined by the spanning vectors $\mathbf{T}_{n}\cdot(\mathbf{a_{1}},\mathbf{a_{2}})$, $n=1,2$, with $\mathbf{T}_{1}=(n_{x_{1}},n_{y_{1}})$ and $\mathbf{T}_{2}=(n_{x_{2}},n_{y_{2}})$ and where $\mathbf{a}_{1},\mathbf{a}_{2}$ are the moiré lattice vectors. The system encloses $N_{s}=|n_{x_{1}}n_{y_{2}}-n_{x_{2}}n_{y_{1}}|$ moiré unit cells [74]. To achieve a uniform sampling of the momentum points in the moiré Brillouin zone, we select the spanning vectors for each system size as follows:
  • $N_{\text{s}}=10$: $\mathbf{T}_{1}=(3,1)$, $\mathbf{T}_{2}=(-1,3)$.
  • $N_{\text{s}}=11$: $\mathbf{T}_{1}=(3,-2)$, $\mathbf{T}_{2}=(1,3)$.
  • $N_{\text{s}}=12$: $\mathbf{T}_{1}=(4,0)$, $\mathbf{T}_{2}=(0,3)$.
  • $N_{\text{s}}=13$: $\mathbf{T}_{1}=(3,-2)$, $\mathbf{T}_{2}=(2,3)$.
  • $N_{\text{s}}=14$: $\mathbf{T}_{1}=(4,2)$, $\mathbf{T}_{2}=(-3,2)$.
  • $N_{\text{s}}=16$: $\mathbf{T}_{1}=(4,0)$, $\mathbf{T}_{2}=(0,4)$.
  • $N_{\text{s}}=18$: $\mathbf{T}_{1}=(4,1)$, $\mathbf{T}_{2}=(-2,4)$.
  • $N_{\text{s}}=20$: $\mathbf{T}_{1}=(5,0)$, $\mathbf{T}_{2}=(0,4)$.
Figure 4: Additional data for Abelian states at half filling with long-range interactions. (a) Many-body energy spectrum (10 lowest states) for $N_{\text{s}}=20$ sites. The two degenerate ground states are marked in red, while the blue arrow denotes the energy gap. (b) Scaling of the energy gap including calculations for $N_{\text{s}}=10,12,14,16,18,20$ sites.
Figure 5: Additional data for non-Abelian states at filling one with long range interaction in a system of $N_{\text{x}}=14$ excitons. (a)-(b) Spectral flow for the momentum sectors $K=1$ and $K=8$, respectively, showing that the ground states remain separated from excited states. (c)-(d) PES for the three ground states with $N_{A}=4$ and $N_{A}=5$, respectively. The number of states below the red line in (c) and (d) is 1848 and 4942, respectively, corresponding to the quasi-hole excitation counting for Moore–Read states. The inset in (d) displays a zoomed-in region in the PES showing the sparse distribution of states where the entanglement gap is expected.
Figure 6: Moore–Read states at filling one for systems with $N_{\text{x}}=11$ and $N_{\text{x}}=12$ excitons. Many-body spectrum (10 lowest energies) and PES at filling $\nu=1$ of the exciton Chern band with long-range interactions for (a)-(b) 11 and (c)-(d) 12 particles. The number of states below the red line matches the quasi-hole counting for the Moore–Read state (253 and 328) for 11 and 12 particles with $N_{A}=3$.
Appendix C Supporting numerical evidence
C.1 Abelian states at half filling
The observed signatures for Laughlin states at half filling remain robust across the wide range of system sizes explored. In Fig. 4(a), we show that the two ground states appear in the energy spectra also for the largest system size that is accessible in our numerical calculations, $N_{\text{s}}=20$. These states are present in smaller systems as well. Importantly, the gap with respect to the excited states remains robust ($\sim 10$ meV) for all the studied systems ranging from $N_{\text{s}}=10$ up to $N_{\text{s}}=20$ sites, cf. Fig. 4(b).
C.2 Non-Abelian states at filling one
In the case of Moore–Read states at filling $1$, the gap between ground and excited states for an even number of excitons ($N_{\text{x}}=14$) is small. Nevertheless, we show in Fig. 5(a)-(b) that ground and excited states remain well separated upon threading a magnetic flux. In addition, in Fig. 5(c)-(d) we show that the presence of a PES gap matching the counting of quasi-hole excitations is a robust feature also for a higher number of particles in the $A$ sub-system, $N_{A}=4$, and is still present for $N_{A}=5$, although less pronounced. In Fig. 6 we show that the characteristic features of Moore–Read states are present not only in the largest systems studied ($N_{\text{x}}=13,14$) but also in systems with a lower number of excitons ($N_{\text{x}}=11,12$ is shown here).